OpenStudy (anonymous):

The manufacturing cost of a cell phone was $124. Sellers first marked up the cost by 20 percent. Sales were poor, so sellers decreased the markup to 15 percent. Calculate the old and new selling prices.

OpenStudy (anonymous):

So what do you know about percentages?

OpenStudy (anonymous):

from 20 decreased to 15

OpenStudy (anonymous):

No, percentages in general. Percentages are the number given over 100. So 20% = 20/100

OpenStudy (juana02):

old prices add $124 by 20%=

OpenStudy (juana02):

new price $124 +15%=?

OpenStudy (anonymous):

right i understand so.. here are my answer choices a- old price 170.00 new price 148. b- old 142.60 new 148.80 c- old 148.80 new 142.60

OpenStudy (anonymous):

So do the calculations... \[(124*20/100)+124= price old\]

OpenStudy (anonymous):

i got 124.15 but not sure if thats right @juana02

OpenStudy (anonymous):

thats wrong

OpenStudy (anonymous):

so 148.80 old and (124*15/100)+124= 142.60 :) hmmm looks good right?

OpenStudy (anonymous):

yup great

OpenStudy (anonymous):

thanks both of u i wish i was able to give medal to more than one person but uhh wtev..thank anyway

OpenStudy (juana02):

thanks @cmvbubu tag me if u have other questions